# On the 2-torsion in class groups of number fields

### v1

@@ Abstract.
In $2020$, Bhargava, Shankar, Taniguchi, Thorne, Tsimerman, and Zhao proved that for a finite extension $K/\mathbb{Q}$ of degree $n\geq 5$, the size of the 22-torsion class group is bounded by $h_{2}(K)=O_{n,\varepsilon}(D_{K}^{\frac{1}{2}-\frac{1}{2n}+\varepsilon})$, where $D_{K}$ is the absolute discriminant of $K$. In the present paper, we improve their bound by proving that $h_{2}(K)=O_{n,\varepsilon}(D_{K}^{\frac{1}{2}-\frac{1}{2n}-\delta_{K}+\varepsilon})$, for a constant $\delta_{K}\geq\frac{1}{28n}-\frac{3}{28n(n-1)}$.
@@ 1. Introduction
where $D_{K}$ denotes the absolute value of the discriminant of $K$ over $\mathbb{Q}$. Recently, Lemke Oliver and Zaman ([LZ25, Corollary 11]) showed that for every $\ell$, one has $h_{\ell}(K)=o(D_{K}^{\frac{1}{2}})$ (and where the saving is an explicit power of $\log(D_{K})$). On the other hand, it is conjectured that
@@ 2. Outline of the proof
Our starting point is the same as [BSTTZ20, Theorem $1.1$]: The size of the 22-torsion of a number field extension is bounded by
@@
That is, when $0\leq\delta_{2,K}\leq\frac{1}{2n}$, Theorem 1.1 holds for $\delta_{K}=\frac{1}{6n}-\frac{\delta_{2,K}}{3}$. Moreover, since $\frac{1}{6n}-\frac{\delta_{2,K}}{3}<0$ when $\delta_{2,K}\geq\frac{1}{2n}$, it follows the bound in Theorem 2.2 is actually worst then the one in [BSTTZ20, Theorem $1.1$] in the range $\frac{1}{2n}\leq\delta_{2,K}\leq\frac{1}{2(n-1)}+o(1)$; hence, Theorem 2.2 alone would not suffice to prove Theorem $1.$
$2.$ The case $\frac{1}{2n}\leq\delta_{2,K}\leq\frac{1}{2(n-1)}+o(1)$.
@@
one finally proves Theorem $1.$
@@
exhibits square root cancellation for all $(r,t)$ which are not contained in a set of codimension 0 in $\mathbb{F}_{p}^{2}$. Then Proposition 2.4 is the analogue of showing that the AA-number associated to the exponential sum 4, in the sense of [Kat89], is not zero. We remark that, in general, determining whether the AA-number associated with an exponential sum does not vanishes is a difficult problem (see [Kat94], [FK01]).
@@ 4.1.1. The contribution of $Z_{\square}$
We start by recording the following Lemma, which is proved in [BSTTZ20, Bottom of page 66] and will be useful in bounding the contribution of the $\operatorname{\mathbf{z}}\in Z_{\square}$:
@@
Assume that $K/\mathbb{Q}$ does not contain a subfield of index 22. Then $Z_{\square}$ is empty if $n=5,6,7$, and
@@
If the polynomial $F_{\operatorname{\mathbf{z}}}(X_{1},X_{2})$ is a square, then $F_{\operatorname{\mathbf{z}}}(X_{1},x_{2})=N_{K\mathbb{Q}}(X_{1}+x_{2}\alpha_{2,K}+z_{3}x_{3}+\cdots+z_{n}\alpha_{n,K})$ is a square for every value of $x_{2}\in\mathbb{Z}$. On the other hand, since we are assuming that $K$ does not contain any subfield of index 22, one can find $(x_{2},\operatorname{\mathbf{z}})$ such that $F_{\operatorname{\mathbf{z}}}(X_{1},x_{2})=N_{K\mathbb{Q}}(X_{1}+x_{2}\alpha_{2,K}+z_{3}x_{3}+\cdots+z_{n}\alpha_{n,K})$ is a square only if there exists $F\subset K$ such that $[K:F]>2$ even (see [BSTTZ20, proof of Lemma 4.1]). Hence, for $n=5,6,7$ the set $Z_{\square}$ is actually empty. Let us assume now $n\geq 8$. We start by writing
@@ 2. $b)$
Corollary $4.$
Let $K/\mathbb{Q}$ be a number field of degree $n\geq 5$ which does not contain any subfield of index 22. Let $\sigma,\tau\in\operatorname{Emb}(K/\mathbb{Q})$ be two distinct element in $\operatorname{Emb}(K/\mathbb{Q})$ such that $\tau(\alpha_{2,K})\neq\alpha_{2,K}$. Then
@@
We first handle the case when $\frac{\sigma(\alpha_{2,K})-\alpha_{2,K}}{\tau(\alpha_{2,K})-\alpha_{2,K}}=1$. In this case, $x$ satisfies $(\tau^{-1}\circ\sigma)(x)=x$, i.e. $x\in K^{\tau^{-1}\circ\sigma}$ which is a $\mathbb{Q}$-vectorspace of dimension at most $\frac{n}{3}$ (since we are assuming that $K$ does not contain any subfield of index 22). Since $1\in V_{\sigma,\tau}$, we have
@@
where $K^{\sigma}$ is the field of all the element in $K$ which are fixed by $\sigma$. Since $K:K^{\sigma}]\geq 3$ (we are assuming that $K$ does not contain any intermediate field of index 22), then $K^{\sigma}$ is a $\mathbb{Q}$-vector field of dimension $\leq\frac{n}{3}$, for every $\sigma\in\operatorname{Emb}(K/\mathbb{Q})\setminus\{\operatorname{Id}\}$. Moreover, since $1\in K^{\sigma}$
@@
Assume that $K/\mathbb{Q}$ does not contain a subfield of index 22, and $\delta_{2,K}\geq\frac{1}{4n}$. Then
@@
Since the bound for $T(\operatorname{\mathbf{u}},\operatorname{\mathbf{z}},pq)$ depends on the G.C.D. between $u_{1},u_{2}$ and pqpq we start by rewriting $S(\operatorname{\mathbf{z}},pq)$ as
@@
We can now prove Theorem $2.$ Thanks to Lemma 4.1, we may assume that $K$ does not contain a subfield of index 22. We recall that
@@
when $n\geq 8$ and zero otherwise. Since $\frac{1}{4}+\frac{1}{n}\leq\frac{1}{2}-\frac{2}{3n}$, for $n>6$, we complete the proof of Theorem $2.$
$4.$ The square sieve for composite moduli: bounding $h_{2}(K)$ when $\delta_{2,K}$ is large.
In this section, we are going to prove Theorem 2.3. As mentioned in the introduction, we do this by deploying the square sieve for composite moduli. We record here the sieve Lemma as stated in [HP12, Lemma 22]
@@
$\sum_{k}\omega_{\operatorname{\mathbf{z}}}(k^{2})\ll_{n,\varepsilon}D_{K}^{\frac{45}{28n}-\frac{11}{14}\delta_{2,K}+\varepsilon}$
@@
provided that $\mathcal{U},\mathcal{V}\subset\{p:F(X,Y)\mod p\text{ is not a square}\}.$ Thus, it remains to estimate $\Sigma_{2B}$. To do this, we first denote $g_{h}(x_{2})=g_{2}(x_{2})g_{2}(x_{2}+q_{2}h)$. Applying the Poisson summation formula, one gets
@@
We can now prove Theorem 2.3. Again, thanks to Lemma 4.1, we may assume that $K$ does not contain a subfield of index 22. We recall that
@@
$\begin{split}\sum_{\begin{subarray}{c}\operatorname{\mathbf{z}}\in\mathcal{C}\cap\mathbb{Z}^{n-2}\\ \operatorname{\mathbf{z}}\not\in Z_{\square}\cup Z_{F}\end{subarray}}\sum_{\begin{subarray}{c}(x_{1},x_{2})\in\mathcal{B}\cap\mathbb{Z}^{2}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1&\ll_{n,\varepsilon}\#(\mathcal{C}\cap\mathbb{Z}^{n-2})\cdot D_{K}^{\frac{45}{28n}-\frac{11}{14}\delta_{2,K}}\\ &\ll D_{K}^{\sum_{j=3}^{n}\left(\frac{1}{n}-\delta_{j,K}\right)+\frac{1}{2}-\frac{1}{2n}-\left(\frac{1}{7n}-\frac{3\delta_{2,K}}{14}\right)+\varepsilon}\\ &=D_{K}^{\frac{1}{2}-\frac{1}{2n}-\left(\frac{1}{7n}-\frac{3\delta_{2,K}}{14}\right)+\varepsilon}.\end{split}$
@@ 5.1. Notions of the formalism of $\ell$-adic trace functions
    The sheaf $\mathcal{F}$ is said to be a middle-extension $\ell$-adic sheaf, punctually pure of weight 0 if it is a middle-extension sheaf and if for every $n\geq 1$ and every $x\in U(\mathbb{F}_{q^{n}})$, the images of the eigenvalues of $(\operatorname{Fr}_{q^{n}}|\mathcal{F}_{\overline{x}})$ via the fixed embedding $\iota:\overline{\mathbb{Q}}_{\ell}\hookrightarrow\mathbb{C}$, are complex numbers of modulus 11.
@@ 3. $iii)$
    For any $x\in\overline{\mathbb{P}}_{\mathbb{F}_{q}^{1}}^{1}(\overline{\mathbb{F}}_{q})$, $\operatorname{Swan}_{x}(j_{*}j^{*}\mathcal{F})$ is the Swan conductor of $\mathcal{F}$ at $x$ (see [Kat88, Chapter 11] for the definition of the Swan conductor).
@@
It turns out that we can see this function as a trace function: combining [FKM19, Definition $5.9$, Theorem $5.10$] and arguing like in [FKM19, Corollary $5.10$] one shows that there exists a middle extension $\ell$-adic sheaf $\mathcal{H}_{F,\chi}$ pure of weight 11 such that
@@ 5.1.1. A 11-dimensional exponential sum
thanks to Theorem $5.$ It rests to prove that $F(X,s^{\prime}X+k^{\prime})$ is square-free for all but $O_{n}(1)$ values of $k^{\prime}$. First observe that
@@
Hence, if we denote by $\chi_{m}$ the unique character of order 22 over $\mathbb{F}_{p^{m}}$, it follows from Lemma 5.8 that
@@
for all but $O_{n}(1)$ values of $k$. Thus, we get the result by applying Theorem $5.$

### v2

@@ Abstract.
In $2020$, Bhargava, Shankar, Taniguchi, Thorne, Tsimerman, and Zhao proved that for a finite extension $K/\mathbb{Q}$ of degree $n\geq 5$, the size of the $2$-torsion class group is bounded by $h_{2}(K)=O_{n,\varepsilon}(D_{K}^{\frac{1}{2}-\frac{1}{2n}+\varepsilon})$, where $D_{K}$ is the absolute discriminant of $K$. In the present paper, we improve their bound by proving that $h_{2}(K)=O_{n,\varepsilon}(D_{K}^{\frac{1}{2}-\frac{1}{2n}-\delta_{K}+\varepsilon})$, for a constant $\delta_{K}\geq\frac{1}{28n}-\frac{3}{28n(n-1)}$.
@@ 1. Introduction
where $D_{K}$ denotes the absolute value of the discriminant of $K$ over $\mathbb{Q}$. Recently, Lemke Oliver and Zaman ([LZ25, Corollary $1$]) showed that for every $\ell$, one has $h_{\ell}(K)=o(D_{K}^{\frac{1}{2}})$ (and where the saving is an explicit power of $\log(D_{K})$). On the other hand, it is conjectured that
@@ 2. Outline of the proof
Our starting point is the same as [BSTTZ20, Theorem $1.1$]: The size of the $2$-torsion of a number field extension is bounded by
@@
That is, when $0\leq\delta_{2,K}\leq\frac{1}{2n}$, Theorem 1.1 holds for $\delta_{K}=\frac{1}{6n}-\frac{\delta_{2,K}}{3}$. Moreover, since $\frac{1}{6n}-\frac{\delta_{2,K}}{3}<0$ when $\delta_{2,K}\geq\frac{1}{2n}$, it follows the bound in Theorem 2.2 is actually worst then the one in [BSTTZ20, Theorem $1.1$] in the range $\frac{1}{2n}\leq\delta_{2,K}\leq\frac{1}{2(n-1)}+o(1)$; hence, Theorem 2.2 alone would not suffice to prove Theorem 1.1.
2.2. The case $\frac{1}{2n}\leq\delta_{2,K}\leq\frac{1}{2(n-1)}+o(1)$.
@@
one finally proves Theorem 1.1.
@@
exhibits square root cancellation for all $(r,t)$ which are not contained in a set of codimension 0 in $\mathbb{F}_{p}^{2}$. Then Proposition 2.4 is the analogue of showing that the $A$-number associated to the exponential sum 4, in the sense of [Kat89], is not zero. We remark that, in general, determining whether the $A$-number associated with an exponential sum does not vanishes is a difficult problem (see [Kat94], [FK01]).
@@ 4.1.1. The contribution of $Z_{\square}$
We start by recording the following Lemma, which is proved in [BSTTZ20, Bottom of page $6$] and will be useful in bounding the contribution of the $\operatorname{\mathbf{z}}\in Z_{\square}$:
@@
Assume that $K/\mathbb{Q}$ does not contain a subfield of index $2$. Then $Z_{\square}$ is empty if $n=5,6,7$, and
@@
If the polynomial $F_{\operatorname{\mathbf{z}}}(X_{1},X_{2})$ is a square, then $F_{\operatorname{\mathbf{z}}}(X_{1},x_{2})=N_{K\mathbb{Q}}(X_{1}+x_{2}\alpha_{2,K}+z_{3}x_{3}+\cdots+z_{n}\alpha_{n,K})$ is a square for every value of $x_{2}\in\mathbb{Z}$. On the other hand, since we are assuming that $K$ does not contain any subfield of index $2$, one can find $(x_{2},\operatorname{\mathbf{z}})$ such that $F_{\operatorname{\mathbf{z}}}(X_{1},x_{2})=N_{K\mathbb{Q}}(X_{1}+x_{2}\alpha_{2,K}+z_{3}x_{3}+\cdots+z_{n}\alpha_{n,K})$ is a square only if there exists $F\subset K$ such that $[K:F]>2$ even (see [BSTTZ20, proof of Lemma 4.1]). Hence, for $n=5,6,7$ the set $Z_{\square}$ is actually empty. Let us assume now $n\geq 8$. We start by writing
@@ 2. $b)$
Corollary 4.4.
Let $K/\mathbb{Q}$ be a number field of degree $n\geq 5$ which does not contain any subfield of index $2$. Let $\sigma,\tau\in\operatorname{Emb}(K/\mathbb{Q})$ be two distinct element in $\operatorname{Emb}(K/\mathbb{Q})$ such that $\tau(\alpha_{2,K})\neq\alpha_{2,K}$. Then
@@
We first handle the case when $\frac{\sigma(\alpha_{2,K})-\alpha_{2,K}}{\tau(\alpha_{2,K})-\alpha_{2,K}}=1$. In this case, $x$ satisfies $(\tau^{-1}\circ\sigma)(x)=x$, i.e. $x\in K^{\tau^{-1}\circ\sigma}$ which is a $\mathbb{Q}$-vectorspace of dimension at most $\frac{n}{3}$ (since we are assuming that $K$ does not contain any subfield of index $2$). Since $1\in V_{\sigma,\tau}$, we have
@@
where $K^{\sigma}$ is the field of all the element in $K$ which are fixed by $\sigma$. Since $K:K^{\sigma}]\geq 3$ (we are assuming that $K$ does not contain any intermediate field of index $2$), then $K^{\sigma}$ is a $\mathbb{Q}$-vector field of dimension $\leq\frac{n}{3}$, for every $\sigma\in\operatorname{Emb}(K/\mathbb{Q})\setminus\{\operatorname{Id}\}$. Moreover, since $1\in K^{\sigma}$
@@
Assume that $K/\mathbb{Q}$ does not contain a subfield of index $2$, and $\delta_{2,K}\geq\frac{1}{4n}$. Then
@@
Since the bound for $T(\operatorname{\mathbf{u}},\operatorname{\mathbf{z}},pq)$ depends on the G.C.D. between $u_{1},u_{2}$ and $pq$ we start by rewriting $S(\operatorname{\mathbf{z}},pq)$ as
@@
We can now prove Theorem 2.2. Thanks to Lemma 4.1, we may assume that $K$ does not contain a subfield of index $2$. We recall that
@@
when $n\geq 8$ and zero otherwise. Since $\frac{1}{4}+\frac{1}{n}\leq\frac{1}{2}-\frac{2}{3n}$, for $n>6$, we complete the proof of Theorem 2.2.
4.4. The square sieve for composite moduli: bounding $h_{2}(K)$ when $\delta_{2,K}$ is large.
In this section, we are going to prove Theorem 2.3. As mentioned in the introduction, we do this by deploying the square sieve for composite moduli. We record here the sieve Lemma as stated in [HP12, Lemma $2$]
@@
$\sum_{k}\omega_{\operatorname{\mathbf{z}}}(k^{2})\ll_{n,\varepsilon}D_{K}^{\frac{2}{n}-\delta_{2,K}-\frac{1}{2n}-\left(\frac{1}{7n}+\frac{3\delta_{2,K}}{14}\right)+\varepsilon}$
@@
provided that $\mathcal{U},\mathcal{V}\subset\{p:F_{\operatorname{\mathbf{z}}}(X,Y)\mod p\text{ is not a square}\}.$ Thus, it remains to estimate $\Sigma_{2B}$. To do this, we first denote $g_{h}(x_{2})=g_{2}(x_{2})g_{2}(x_{2}+q_{2}h)$. Applying the Poisson summation formula, one gets
@@
We can now prove Theorem 2.3. Again, thanks to Lemma 4.1, we may assume that $K$ does not contain a subfield of index $2$. We recall that
@@
$\begin{split}\sum_{\begin{subarray}{c}\operatorname{\mathbf{z}}\in\mathcal{C}\cap\mathbb{Z}^{n-2}\\ \operatorname{\mathbf{z}}\not\in Z_{\square}\cup Z_{F}\end{subarray}}\sum_{\begin{subarray}{c}(x_{1},x_{2})\in\mathcal{B}\cap\mathbb{Z}^{2}\\ F_{\operatorname{\mathbf{z}}}(x_{1},x_{2})=\square\end{subarray}}1&\ll_{n,\varepsilon}\#(\mathcal{C}\cap\mathbb{Z}^{n-2})\cdot D_{K}^{\frac{2}{n}-\delta_{2,K}-\frac{1}{2n}-\left(\frac{1}{7n}+\frac{3\delta_{2,K}}{14}\right)+\varepsilon}\\ &\ll D_{K}^{\sum_{j=3}^{n}\left(\frac{1}{n}-\delta_{j,K}\right)+\frac{2}{n}-\delta_{2,K}-\frac{1}{2n}-\left(\frac{1}{7n}+\frac{3\delta_{2,K}}{14}\right)+\varepsilon}\\ &=D_{K}^{\frac{1}{2}-\frac{1}{2n}-\left(\frac{1}{7n}-\frac{3\delta_{2,K}}{14}\right)+\varepsilon}.\end{split}$
@@ 5.1. Notions of the formalism of $\ell$-adic trace functions
    The sheaf $\mathcal{F}$ is said to be a middle-extension $\ell$-adic sheaf, punctually pure of weight 0 if it is a middle-extension sheaf and if for every $n\geq 1$ and every $x\in U(\mathbb{F}_{q^{n}})$, the images of the eigenvalues of $(\operatorname{Fr}_{q^{n}}|\mathcal{F}_{\overline{x}})$ via the fixed embedding $\iota:\overline{\mathbb{Q}}_{\ell}\hookrightarrow\mathbb{C}$, are complex numbers of modulus $1$.
@@ 3. $iii)$
    For any $x\in\overline{\mathbb{P}}_{\mathbb{F}_{q}^{1}}^{1}(\overline{\mathbb{F}}_{q})$, $\operatorname{Swan}_{x}(j_{*}j^{*}\mathcal{F})$ is the Swan conductor of $\mathcal{F}$ at $x$ (see [Kat88, Chapter $1$] for the definition of the Swan conductor).
@@
It turns out that we can see this function as a trace function: combining [FKM19, Definition $5.9$, Theorem $5.10$] and arguing like in [FKM19, Corollary $5.10$] one shows that there exists a middle extension $\ell$-adic sheaf $\mathcal{H}_{F,\chi}$ pure of weight $1$ such that
@@ 5.1.1. A $1$-dimensional exponential sum
thanks to Theorem 5.5. It rests to prove that $F(X,s^{\prime}X+k^{\prime})$ is square-free for all but $O_{n}(1)$ values of $k^{\prime}$. First observe that
@@
Hence, if we denote by $\chi_{m}$ the unique character of order $2$ over $\mathbb{F}_{p^{m}}$, it follows from Lemma 5.8 that
@@
for all but $O_{n}(1)$ values of $k$. Thus, we get the result by applying Theorem 5.5.